# Cobham’s Theorem for the Gaussian integers

### v1

@@ 1. Introduction
A sequence $a=(a_{n})_{n\geq 0}$ is $k$-automatic if $a_{n}$ is a finite-state function of the base-$k$ expansion of $n$. Let $k$ and $\ell$ be multiplicatively independent positive integers. Cobham’s theorem [9] tells us that a sequence is both $k$- and $\ell$-automatic if and only if it is eventually periodic. This theorem is arguably the most significant result concerning automatic sequences. Its importance is reflected in the fact that it is connected to many deep generalisations and proofs in different areas of mathematics, such as the work of Semenov and Muchnik in logic [24, 20], the work of Durand in dynamical systems and substitution theory [11, 12], the work of Bès in numeration [4], the work of Adamczewski and Bell in algebra [3, 1], and the work of the latter and also Schäfke and Singer in analysis and ODEs [2, 23]. This list is far from exhaustive.
While Cobham’s theorem has been extended in many directions, one notable gap is its extension from the integers to Euclidean domains. In this paper, we take a first step and extend Cobham’s theorem to the Gaussian integers. To formulate the theorem in this context, however, one must first establish what a $\beta$-automatic configuration is for a Gaussian integer $\beta$. In particular, this requires a numeration system with base $\beta$. The numeration system for the Gaussian integers with base $\beta=-1+i$ and digits 0,1 is described by Knuth [16, Section 4.1], who attributes it to Penney[22]. The only Gaussian integers for which any $z\in\mathbb{Z}[\mathrm{i}]$ has an expansion base-$\alpha$ with natural numbers as digits, are Gaussian integers $\alpha=-m\pm\mathrm{i}$, with digit set $\{0,1,\dots,m^{2}\}$, as shown by Kátai and Szabó in [15].
In [14], Hansel and Safer considered these Gaussian integers $\alpha=-m+\mathrm{i}$ as a base for Gaussian numeration systems ($m\geq 1$), with digit set $\{0,1,\dots,m^{2}\}$. Any two Gaussian integers of this form must be multiplicatively independent. Assuming the four exponentials conjecture, Hansel and Safer showed that for such Gaussian integers, if the characteristic function of an infinite $S\subset\mathbb{Z}[i]$ is both $\alpha$- and $\beta$-recognizable, then it is syndetic. They conjectured that this set must be eventually periodic, where we recall that a configuration $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ is eventually periodic if there exist $\mathbb{Z}$-linearly independent $p,q\in\mathbb{Z}[\mathrm{i}]$ such that $a_{z}=a_{z+p}=a_{z+q}$ whenever z[$\Fz\in\mathbb{Z}[\mathrm{i}]\backslash$ F for some finite set $F$.
@@
Let $\alpha$ and $\beta$ be two multiplicatively independent Gaussian integers with $|\alpha|,|\beta|>1$. If one of $\alpha$ or $\beta$ is not the root of an integer, then a configuration is $\alpha$- and $\beta$-automatic if and only if it is eventually periodic.
In particular, Hansel and Safer’s conjecture is proved true, since m+i-m+i is not the root of an integer if $m>1$. The requirement that one of $\alpha$ and $\beta$ must be a root of an integer cannot be relaxed, as the characteristic function of $\mathbb{N}$ can be shown to be $\alpha$-automatic whenever $\alpha$ is a root of an integer. Multiplicatively independent Gaussian roots of integers will however obey a Semenov-type theorem [24, 20], see also the exposition [6], behaving like higher dimensional automatic configurations generated by a base-$k$ numeration system for $\mathbb{N}\times\mathbb{N}$, where projections of $\alpha$- and $\beta$-automatic configurations are eventually periodic along any horizontal or vertical direction.
@@ 2.1. Gaussian numerations
We assume the reader is familiar with the fact that the Gaussian integers form a Euclidean domain. In what follows, we discuss some basic facts about the Gaussian integers and their numeration systems that shall be of use to us. For a more in-depth study of Gaussian numeration systems, we direct the reader to [18, Chapter 7].
For $z\in\mathbb{Z}[\mathrm{i}]$, we let $|z|$ denote the Euclidean norm of $z$. Given a Gaussian integer $\gamma\in\mathbb{Z}[\mathrm{i}]$ with $\lvert\gamma\rvert>1$ and a finite subset $D\subset\mathbb{Z}[\mathrm{i}]$ containing 0, to each word $w=w_{n-1}\dotsc w_{0}\in D^{*}$ we associate the number
@@
note that $w_{0}$ is the least significant digit. This notation extends to sets of words $L\subseteq D^{*}$, by writing $L]_{\gamma}=\{[w]_{\gamma}\>:\>w\in L\}$. We say that $w\in D^{*}$ is a $(\gamma,D)$-expansion of $z\in\mathbb{Z}[\mathrm{i}]$ if $w]_{\gamma}=z$. We say that $(\gamma,D)$ is an integral numeration system if every Gaussian integer $z$ has a unique expansion $n=\sum_{j=0}^{k_{n}-1}d_{j}\gamma^{j}$ with $d_{j}\in D$. In this case we write $(n)_{\gamma}=d_{k}\dots d_{0}$. A set $D\in\mathbb{Z}[\mathrm{i}]$ is a complete residue system for $\mathbb{Z}[\mathrm{i}]\pmod{\gamma}$ if every Gaussian $z$ is congruent modulo $\gamma$ to a unique element of $D$. We record the following result [18, Theorems 7.4.1, 7.4.2, 7.4.3], which indicates that we have an abundance of Gaussian numeration systems.
@@ 2.2. Dirichlet approximation in $\mathbb{Z}[\mathrm{i}]$
In his original proof, Cobham needed the fact that for natural numbers $\alpha,\beta\geq 2$, the multiplicative group generated by the natural $\alpha$ and $\beta$ is dense in the positive real line. It turns out that this result is very hard to generalise to $\mathbb{C}$. Hansel and Safer showed that if $\alpha$ and $\beta$ are multiplicatively independent Gaussians, then the four exponentials conjecture implies that the multiplicative group $\langle\alpha,\beta\rangle^{\times}=\{\alpha^{m}\beta^{n}\>:\>m,n\in\mathbb{Z}\}$ is dense in $\mathbb{C}$. In our proof of Theorem 1, we only need that there are no isolated points in $\langle\alpha,\beta\rangle^{\times}$, a much weaker result.
@@
Hansel and Safer considered basis $\alpha=-m+i$ for $m\geq 1$. It turns out that the only numbers that are multiplicatively dependent on $\alpha$ are powers of $\alpha$.
@@
Let $y=N(\beta)$, the norm of $\beta$. Taking norms we find $y^{j}=(m^{2}+1)^{k}$. Therefore, $j\cdot\text{ord}_{p}(y)=k\cdot\text{ord}_{p}(m^{2}+1)$ for all (rational) primes $p$. By dividing out common divisors, we may suppose that $\text{gcd}(j,k)=1$. It follows that $k$ divides all $\text{ord}_{p}(y)$. In particular, $y=z^{k}$ for some $z$ such that $z^{j}=m^{2}+1$. By Mihăilescu’s theorem [19], $j=1$. Since we divided out a common factor $h$, it follows that $\beta^{h}=\alpha^{hk}$. Up to a root of unity, $\beta$ is equal to $\alpha^{k}$. ∎
@@ 2.3. Finite automata and automaticity
We will need to consider automatic configurations defined by DFAOs with redundant digit sets, i.e. Gaussian numerations where each Gaussian integer has at least one expansion.
@@
The following is based on [18, Lemma 7.1.1]; we include it for completeness.
@@
The idea is as follows. We go through each digit of an expansion $d_{n}^{\prime}\dotsc d_{0}^{\prime}\in(D^{\prime})^{*}$ for a Gaussian integer $z\in\mathbb{Z}[\mathrm{i}]$, starting from the least significant digit $d_{0}^{\prime}$, which we replace by the unique digit $d_{0}\in D$ which is congruent to $d_{0}^{\prime}$ modulo $\beta$. We now have a carry $c_{0}$ for which the equality $\beta c_{0}=d_{0}^{\prime}-d_{0}$ holds, so $z=[d_{n}^{\prime}\cdots d_{1}^{\prime}d_{0}]_{\beta}+\beta c_{0}$. We iterate this process, at each step defining $c_{k}=d_{k}^{\prime}-d_{k}$ and $d_{k+1}$ as the only element of $D$ congruent to $c_{k}+d_{k+1}^{\prime}$, so that $d_{k+1}^{\prime}+c_{k}=\beta c_{k+1}+d_{k+1}$ for some value $c_{k+1}\in\mathbb{Z}[\mathrm{i}]$, which will become the next carry. At each step the equality $z=[d_{n}^{\prime}\dotsc d_{k+1}^{\prime}d_{k}\dotsc d_{0}]_{\beta}+\beta^{k}c_{k}$ holds. After we modify all of the digits $d_{0}^{\prime},\dotsc,d_{n}^{\prime}$, we have an expression of the form $z=\beta^{n}c_{n}+[d_{n}\dotsc d_{0}]_{\beta}$, so we only need to append the digits of the $(\beta,D)$-expansion of $c_{n}$ at the start to get a full $(\beta,D)$-expansion of $z$.
@@
The following proposition is reminiscent of Cabezas and Leroy’s [7, Proposition 4.1], which gives a similar result in the primitive, aperiodic, substitutional case.
@@ 3. Generating periods
Let $\gamma$ be a Gaussian integer with $\lvert\gamma\rvert>1$, and $D$ a digit set for which each $z\in\mathbb{Z}[\mathrm{i}]$ has at least one $(\gamma,D)$-expansion. Suppose that $\mathcal{A}_{\gamma}=(S,D,\delta,s_{0},A,\tau)$ is a consistent $(\gamma,D)$-automaton, and let $s\in S$; we define:
@@ 5. Proof of the main theorem for $\mathbb{Z}[\mathrm{i}]$
The statement of Theorem 1 is the best possible. For, suppose that $\alpha$ and $\beta$ are both roots of an integer. Then there exist configurations which are both $\alpha$- and $\beta$- automatic, but not eventually periodic. To see this we apply [5, Theorem 4], which tells us that $\mathbb{N}$ is a $\beta$-automatic set if and only if $\beta$ is a root of an integer. The characteristic function of $\mathbb{N}$ is not eventually periodic. It can be shown that a Semenov-Cobham type theorem exists for such bases.
@@
Suppose that $U$ is $(\beta,D)$-automatic. If needed, using Proposition 10, we can enlarge $D$ so that it is closed under conjugation. Since $U=\overline{U}$, we conclude that $U$ is $(\bar{\beta},D)$-automatic. Since $U$ is not eventually periodic, $\beta$ and $\bar{\beta}$ are multiplicatively dependent. Thus there exist $m,n$ such that $|\beta|^{n}=|\bar{\beta}|^{m}$. This implies that we can take $m=n$, since $|\beta|>1$. Thus $\beta^{n}=\bar{\beta}^{n}$, so that $\beta^{n}\in\mathbb{R}$. ∎

### v2

@@ 1. Introduction
A sequence $a=(a_{n})_{n\geq 0}$ is $k$-automatic if $a_{n}$ is the output of a function computed by a finite state automaton, taking as input the base-$k$ expansion of $n$. Such sequences are the simplest sequences beyond the eventually periodic sequences. The base of the numeration system matters: Let $k$ and $\ell$ be multiplicatively independent positive integers, then Cobham’s theorem [11] tells us that a sequence is both $k$- and $\ell$-automatic if and only if it is eventually periodic. This theorem is arguably the most significant result concerning automatic sequences; its importance is reflected in the fact that it is connected to many deep generalisations and proofs in different areas of mathematics, such as the work of Semenov and Muchnik in logic [28, 24], the work of Durand and others in dynamical systems and substitution theory [14, 15, 12, 18], the work of Bès in numeration [6], the work of Adamczewski and Bell in algebra [5, 1], and the work of the latter and also Schäfke and Singer in analysis and ODEs [2, 27]. There are also recent exciting algebraic generalisations of Cobham’s theorem by Adamczewski and Faverjon [4, 3].
While Cobham’s theorem has been extended in many directions, one notable gap is its extension from the integers to Euclidean domains. In this paper, we take a first step and extend Cobham’s theorem to the Gaussian integers. To formulate the theorem in this context, however, one must first establish what a $\beta$-automatic configuration is for a Gaussian integer $\beta$. In particular, this requires a numeration system with base $\beta$. The numeration system for the Gaussian integers with base $\beta=-1+i$ and digits 0,1 is described by Knuth [20, Section 4.1], who attributes it to Penney[26]. The only Gaussian integers for which any $z\in\mathbb{Z}[\mathrm{i}]$ has an expansion base-$\alpha$ with natural numbers as digits, are Gaussian integers $\alpha=-m\pm\mathrm{i}$, with digit set $\{0,1,\dots,m^{2}\}$, as shown by Kátai and Szabó in [19].
In [17], Hansel and Safer considered these Gaussian integers $\alpha=-m+\mathrm{i}$ as a base for Gaussian numeration systems ($m\geq 1$), with nonnegative integer digit sets; Any two such Gaussian integers must be multiplicatively independent (see, eg, Theorem 6). Assuming the four exponentials conjecture, they showed that for such Gaussian integers, if the characteristic function $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ of an infinite $S\subset\mathbb{Z}[i]$ is both $\alpha$- and $\beta$-recognizable, then it is syndetic. They conjectured that this set must be eventually periodic, i.e., there exist $\mathbb{Z}$-linearly independent $p,q\in\mathbb{Z}[\mathrm{i}]$ such that $a_{z}=a_{z+p}=a_{z+q}$ whenever z[$\Fz\in\mathbb{Z}[\mathrm{i}]\backslash$ F for some finite set $F$.
@@
Let $\alpha$ and $\beta$ be two multiplicatively independent Gaussian integers with $|\alpha|,|\beta|>1$. If one of $\alpha$ or $\beta$ is not the root of an integer, then a configuration $(a_{z})_{z\in\mathbb{Z}[i]}$ is $\alpha$- and $\beta$-automatic if and only if it is eventually periodic.
Niven’s theorem limits Gaussian integers that are roots of integers to have an argument that is a multiple of $\pi/4$ (Theorem 5). In particular, Hansel and Safer’s conjecture is proved true, since m+i-m+i is not the root of an integer if $m>1$. The requirement that one of $\alpha$ and $\beta$ must be a root of an integer cannot be relaxed, as the characteristic function of $\mathbb{N}$ can be shown to be $\alpha$-automatic whenever $\alpha$ is a root of an integer; see Remark 22. Multiplicatively independent Gaussian roots of integers will however obey a Semenov-type theorem [28, 24], see also the exposition [8], behaving like higher dimensional automatic configurations generated by a base-$k$ numeration system for $\mathbb{N}\times\mathbb{N}$, where projections of $\alpha$- and $\beta$-automatic configurations are eventually periodic along any horizontal or vertical direction.
@@ 2.1. Gaussian numerations
In what follows, we discuss some basic facts about the Euclidean domain of Gaussian integers and their numeration systems. (For a more in-depth study of Gaussian numeration systems, we direct the reader to [22, Chapter 7]). For $z\in\mathbb{Z}[\mathrm{i}]$, we let $|z|$ denote the Euclidean norm of $z$. Given a Gaussian integer $\gamma$ with $\lvert\gamma\rvert>1$ and a finite subset $D\subset\mathbb{Z}[\mathrm{i}]$ containing 0, to each word $w=w_{n-1}\dotsc w_{0}\in D^{*}$ we associate the number
@@
note that $w_{0}$ is the least significant digit. For $L\subseteq D^{*}$ we write $L]_{\gamma}=\{[w]_{\gamma}\>:\>w\in L\}$. We say that $w\in D^{*}$ is a $(\gamma,D)$-expansion of $z\in\mathbb{Z}[\mathrm{i}]$ if $w]_{\gamma}=z$. We say that $(\gamma,D)$ is an integral numeration system if every Gaussian integer $z$ has a unique expansion $n=\sum_{j=0}^{k_{n}-1}d_{j}\gamma^{j}$ with $d_{j}\in D$ and $d_{k_{n}-1}\neq 0$. In this case we write $(n)_{\gamma}=d_{k}\dots d_{0}$. A set $D\in\mathbb{Z}[\mathrm{i}]$ is a complete residue system for $\mathbb{Z}[\mathrm{i}]\pmod{\gamma}$ if every Gaussian $z$ is congruent modulo $\gamma$ to a unique element of $D$. We record the following result [22, Theorems 7.4.1, 7.4.2, 7.4.3], which indicates that we have an abundance of Gaussian numeration systems.
@@ 2.2. Dirichlet approximation in $\mathbb{Z}[\mathrm{i}]$
In his original proof, Cobham needed the fact that for natural numbers $\alpha,\beta\geq 2$, the multiplicative group generated by $\alpha$ and $\beta$ is dense in the positive real line. It turns out that this result is very hard to generalise to $\mathbb{C}$. Hansel and Safer showed that if $\alpha$ and $\beta$ are multiplicatively independent Gaussians, then the four exponentials conjecture implies that the multiplicative group $\langle\alpha,\beta\rangle^{\times}=\{\alpha^{m}\beta^{n}\>:\>m,n\in\mathbb{Z}\}$ is dense in $\mathbb{C}$. In our proof of Theorem 1, we only need that there are no isolated points in $\langle\alpha,\beta\rangle^{\times}$, a much weaker result.
@@
Hansel and Safer considered bases $\alpha=-m+i$ for $m\geq 1$. It turns out that the only numbers that are multiplicatively dependent on $\alpha$ are powers of $\alpha$.
@@
Let $y=N(\beta)$, the norm of $\beta$. Taking norms we find $y^{j}=(m^{2}+1)^{k}$. Therefore, $j\cdot\text{ord}_{p}(y)=k\cdot\text{ord}_{p}(m^{2}+1)$ for all (rational) primes $p$. By dividing out common divisors, we may suppose that $\text{gcd}(j,k)=1$. It follows that $k$ divides $\text{ord}_{p}(y)$. In particular, $y=z^{k}$ for some $z$ such that $z^{j}=m^{2}+1$. By Mihăilescu’s theorem [23], $j=1$. Since we divided out a common factor $h$, it follows that $\beta^{h}=\alpha^{hk}$. Up to a root of unity, $\beta$ is equal to $\alpha^{k}$. ∎
@@ 2.3. Finite automata and automaticity
We will need to consider automatic configurations defined by DFAOs with redundant digit sets, i.e. Gaussian numerations where each Gaussian integer has at least one expansion with nonzero leading digit.
@@
The following is based on [22, Lemma 7.1.1]; we include it for completeness.
@@
The idea is as follows. We go through each digit of an expansion $d_{n}^{\prime}\dotsc d_{0}^{\prime}\in(D^{\prime})^{*}$ for a Gaussian integer $z\in\mathbb{Z}[\mathrm{i}]$, starting from the least significant digit $d_{0}^{\prime}$, which we replace by the unique digit $d_{0}\in D$ which is congruent to $d_{0}^{\prime}$ modulo $\beta$. We now have a carry $c_{0}$ for which the equality $\beta c_{0}=d_{0}^{\prime}-d_{0}$ holds, so $z=[d_{n}^{\prime}\cdots d_{1}^{\prime}d_{0}]_{\beta}+\beta c_{0}$. We iterate this process, at each step defining $c_{k}=(d_{k}^{\prime}-d_{k})/\beta$ and $d_{k+1}$ as the only element of $D$ congruent to $c_{k}+d_{k+1}^{\prime}$, so that $d_{k+1}^{\prime}+c_{k}=\beta c_{k+1}+d_{k+1}$ for some value $c_{k+1}\in\mathbb{Z}[\mathrm{i}]$, which will become the next carry. At each step the equality $z=[d_{n}^{\prime}\dotsc d_{k+1}^{\prime}d_{k}\dotsc d_{0}]_{\beta}+\beta^{k}c_{k}$ holds. After we modify all of the digits $d_{0}^{\prime},\dotsc,d_{n}^{\prime}$, we have an expression of the form $z=\beta^{n}c_{n}+[d_{n}\dotsc d_{0}]_{\beta}$, so we only need to append the digits of the $(\beta,D)$-expansion of $c_{n}$ at the start to get a full $(\beta,D)$-expansion of $z$.
@@
The following proposition is reminiscent of Cabezas and Leroy’s [9, Proposition 4.1], which gives a similar result in the primitive, aperiodic, substitutional case.
@@ 3. Generating periods
Let $\gamma$ be a Gaussian integer with $\lvert\gamma\rvert>1$, and let $D$ be a digit set for which each $z\in\mathbb{Z}[\mathrm{i}]$ has at least one $(\gamma,D)$-expansion. Suppose that $\mathcal{A}_{\gamma}=(S,D,\delta,s_{0},A,\tau)$ is a consistent $(\gamma,D)$-automaton, and let $s\in S$; we define:
@@ 5. Proof of the main theorem for $\mathbb{Z}[\mathrm{i}]$
The statement of Theorem 1 is the best possible. For, suppose that $\alpha$ and $\beta$ are both roots of an integer. Then there exist configurations which are both $\alpha$- and $\beta$- automatic, but not eventually periodic. To see this we apply [7, Theorem 4], which tells us that $\mathbb{N}$ is a $\beta$-automatic set if and only if $\beta$ is a root of an integer. The characteristic function of $\mathbb{N}$ is not eventually periodic. It can be shown that a Semenov-Cobham type theorem exists for such bases.
@@
Suppose that $U$ is $(\beta,D)$-automatic. If needed, using Proposition 10, we can enlarge $D$ so that it is closed under conjugation. Since $U=\overline{U}$, we conclude that $U$ is $(\bar{\beta},D)$-automatic. Since $U$ is not eventually periodic, $\beta$ and $\bar{\beta}$ are multiplicatively dependent. Thus there exist $m,n$ such that $\beta^{n}=\bar{\beta}^{m}$ and so $|\beta|^{n}=|\bar{\beta}|^{m}$. This implies that $m=n$, since $|\beta|>1$. Thus $\beta^{n}=\bar{\beta}^{n}$, so that $\beta^{n}\in\mathbb{R}$. ∎